It says in the book I'm supposed to write an equation and use it to solve the problem. The questions states: five sixths of a number is -60. What is the number. I think the answer is -300 but I don't really know how to show the work correctly or if that is even the right answer? Any thoughts? Thanks.
Answer by rajagopalan(174)

let the number be=x
5/6ths of the number is=(5/6)x
(5/6)x=-60
x=(-60)*6/5
x=-360/5=-72
the number is =-72
verification;
the given number=-72
5/6 of the number=(-72*5)/6=-360/6=-60
okay
good day